plshelpmeoml:

PLSSSSS HELP MY LAST QUESTION What is the value of h when the function is converted to vertex form? Note: Vertex form is g(x)=a(x−h)^2+k . g(x)=x^2−6x+14 Enter your answer in the box. h =

MoonShadow:

g(x) = x2 - 6x + 14 Complete the square: Take ½ of -6 and square that number = x2 - 6x + 9 + 5 = (x-3)2 + 5 So, h = 3 and k = 5 (Vertex = (h,k) = (3,5))

MoonShadow:

Ur welcome

plshelpmeoml:

CORRECT! THANK YOU SM

MoonShadow:

NP